# Convergent transformations of visual representation in brains and models

### v1

@@ 2.2 Model–Brain Alignment Mirrors Inter-Subject Shared Geometry
Figure 3: Hierarchical convergence between models and cortex. (A–C) Layer-wise alignment (RSA) between vision models and brain activity for representative parcels within the three reference hubs. The alignment curves are averaged across all vision models used in the study. Panels show Early Visual Cortex (A), the Ventral Hub (B), and the LOTC Hub (C). Lines show the mean alignment across participants ($N=8$), with shaded areas representing the standard error of the mean (SEM). Alignment profiles differ by hub, showing early, distributed, and late-stage peaks, respectively. (D) Cortical surface map of the vision model layer with the highest alignment for each parcel. The color scale indicates the normalized depth of the peak layer (0% shallower layer, 100% deeper layer), revealing a posterior-to-anterior gradient from shallow (green) to deep (purple) layers. (E) Layer-wise alignment between language model and representative parcels from each of the three hubs. In contrast to vision models, only parcels within the LOTC hub show strong alignment, which follows a step-like function rather than a smooth progression. (F) Box plots showing the distribution of peak alignment depths for the 20 parcels with the highest overall vision-model alignment. Parcels are sorted by their median peak depth, illustrating a continuous hierarchical organization spanning from early visual areas to the LOTC hub.
@@ 2.3 Hierarchical correspondence Between Models and Cortex
To obtain a more precise mapping of how brain regions correspond to different levels of model abstraction, we moved beyond single summary score and examined the layer-wise alignment profiles for each cortical parcel. While numerous studies have established a coarse hierarchical correspondence showing that early model layers align with V1–V4 and deeper layers map onto downstream regions 15, 16, 17, 51, 52, it remains unclear whether this is a simple one-to-one mapping. Alternatively, certain regions might exhibit a mixed alignment by integrating features from multiple computational stages. To address this, we conducted a layer-wise RSA for each parcel, yielding an alignment profile across model depth (see Methods; Fig. 3).
This layer-wise analysis of vision models showed distinct alignment profiles for each of the three main cortical hubs (Fig.3a–c). Alignment with the early visual cortex peaked in the shallowest model layers before decreasing, consistent with its role in processing low-level features (Fig.3a). In contrast, the ventral hub exhibited a broad plateau of moderate alignment across many layers, suggesting it integrates both visual features from earlier stages and more abstract information from deeper stages (Fig.3b). The LOTC hub, in turn, showed a progressively increasing alignment that peaked only in the deepest, most semantic layers (Fig.3c; see Supplementary Figs. D3D4 for extended parcels and model-family details).
To make the functional hierarchy explicit, we mapped the peak alignment layer of vision models for each parcel. Despite the mixed correspondence observed in some ventral regions, this analysis revealed a smooth posterior-to-anterior gradient across the visual system (Fig. 3d). This gradient was most apparent within visual cortex, whereas frontal areas exhibited uniformly flattened alignment profiles, reflecting mixed correspondence with features spanning multiple model stages (see Supplementary Fig. D3). Nevertheless, the ordered progression from low-level to high-level representations was highly consistent across participants in visual areas (Fig.3f), providing evidence for a shared progression toward abstraction mirrored between cortex and models.
Language models, in contrast, showed no cortex-wide progression; strong alignment emerged only in the LOTC hub (Fig. 3e). Elsewhere, values hovered near zero across layers (Supplementary Fig. D3). Within LOTC, the alignment curve followed a step-like profile: it began with a weak score that was highly correlated with the token-level embeddings fed into the first layer (Pearson’s $r=0.78$; tokenizer control, Supplementary Fig. A4), consistent with work that suggest that brain–language alignment may reflect discrete word properties rather than compositional meaning 53. After this initial stage, alignment increased sharply before reaching a stable maximum, with the final value proportional to the first-layer alignment (Pearson’s $r=0.97$). This pattern indicates that LOTC representations are not late-stage vision, but are instead organized around semantic dimensions driven by token representations in language, with alignment increasing once language models have formed stable embedding representations.
Although these analyses showed that some parcels exhibited mixed correspondence with features from different model stages, they revealed a cortical processing hierarchy that was partially mirrored in the architecture of deep vision models. The distinct, non-hierarchical profile observed for language models further highlighted this structure, isolating the LOTC hub as a region organized around a different, non-visual representational principle. Having established the vision model hierarchy as a proxy for the brain’s own computational stages, we proceeded to examine how information is routed between cortical regions.
@@
Figure 4: Inter-subject representational connectivity network. (A) Inter-subject connectivity matrix. Parcel-wise representational connectivity (RSA, Pearson’s $r$) is shown for 30 key cortical parcels ($N=8$). Each cell represents the mean RSA score between the RDMs of two parcels, computed across all pairs of different individuals. The matrix reveals a clear block structure corresponding to three principal hubs: Early Visual Cortex (V1–V4), the Ventral Hub (PHA1-3, VMV1–3), and the LOTC Hub (MT, MST, V4t, FST, TPOJ2–3). Warm colours denote shared representational geometry; cool colours denote dissimilarity. (B) Directed connectivity graph. The same 30 parcels are plotted as nodes. Node colour indicates each parcel’s peak alignment depth with vision models, node size reflects its inter-subject alignment strength, and the node border colour marks its macro-anatomical group. The edges shown are the union of the three minimum-spanning trees (see Methods). Edge colour encodes the RSA value from panel (a), while edge width indicates the spanning tree order (from first to third), highlighting the network’s most central pathways. Arrowheads denote the putative direction of information flow, inferred from significant differences in peak-depth between connected parcels (paired $t$-test, $t(7)$, FDR-corrected $p<0.05$). The layout reveals two primary processing streams emerging from early visual cortex: a medial-ventral route and a lateral-dorsal route projected to the LOTC hub.
2.4 Representational connectivity reveals an ordered processing network
To characterize the large-scale organization of these functionally distinct hubs, we extended our analysis from isolated regional analysis to the connectivity level. We constructed a parcel-by-parcel representational connectivity matrix by correlating the representational geometry of each parcel in one individual with that of every other parcel in all other individuals. This inter-subject representational connectivity approach directly captures a network structure that is reproducibly shared across the population. To estimate the direction of information flow within this network, we used the peak alignment depth from our vision model analysis as a proxy for hierarchical position, under the assumption that information flows from regions preferring shallower layers to those preferring deeper ones–a principle consistent with feed-forward processing in both the visual system and deep networks 16, 17. The same network structure was obtained using a conventional within-subject approach, confirming the robustness of our findings (see Supplementary Fig. D5).
The parcel-by-parcel connectivity matrix revealed three internally connected clusters corresponding to the Early Visual, Ventral, and LOTC hubs (Fig. 4a). The Early Visual and Ventral hubs were moderately connected, forming an integrated system for visually grounded representation. In contrast, the LOTC hub exhibited a dissimilar representational geometry (negative RSA) relative to both the Ventral hub and classical dorsal regions (e.g., intraparietal sulcus). This pattern supports its functional dissociation as a distinct processing pathway, not only from ventral scene-perception areas like the PPA 54, 55, but also from other dorsal regions involved in spatial processing—a dissociation consistent with proposals for a distinct lateral or “third” visual pathway 30, 56.
The directed graph, which highlights the network’s main connections, made the two-stream topology of information flow explicit (Fig. 4b). Representational similarity originated in early visual cortex and bifurcated into two primary branches: a medial-ventral stream that passed through the Ventral hub, and a lateral-dorsal stream that traversed the LOTC hub—mirroring the dual-network organization previously described for scene processing 54. Our whole-cortex rendering (Supplementary Fig. D6) showed these streams extending anteriorly into posterior parietal, prefrontal, and default-mode regions, outlining a plausible route by which perceptual information can reach circuits for attention, memory, decision-making, and motor planning, consistent with known anatomical and functional pathways.
In sum, our representational connectivity analysis identifies a stimulus-driven network whose core structure is consistent across individuals and is consistent with the dual-network topology previously reported for scene perception 54. This network originates in early visual cortex and bifurcates into two primary processing streams. The medial-ventral stream, terminating in the Ventral hub, corresponds to the classical pathway for scene and object recognition 26, 27. The lateral-dorsal stream, culminating in the LOTC hub, is highly consistent with recent proposals for a third visual pathway specialized for social perception 30. This functional dissociation raises the question of what information defines each of these processing routes—a question we address in the next section by isolating the dominant representational dimensions within each hub.
2.5 Content Decomposition of Representational Hubs
Figure 5: Dominant representational dimensions in cortical hubs. (A–C) Stimuli projection onto the first two shared components for each hub, extracted via Kernel Multi-view CCA (KMCCA) using the voxel data from all eight NSD participants for the common image set. In Early Visual Cortex (A), stimuli form a category-free cloud reflecting low-level visual similarity. In the Ventral Hub (B), the dominant axis arranges stimuli along a scene-to-object gradient. In the LOTC Hub (C), the first component separates animate from inanimate stimuli. (D) Impact on inter-subject RSA of controlling for the first KMCCA component via partial RSA. Controlling for the first component significantly reduced alignment in all hubs: Early Visual (6.6%), Ventral (3.1%), and LOTC (21.8%). The pronounced RSA reduction in LOTC underscores the animacy axis as its dominant organizing principle. (E, F) Cortical maps of inter-subject RSA recomputed separately for those scenes without (E) versus with (F) biological agents (people or animals). Alignment in the LOTC hub collapses when biological content is absent. Partial RSA values in (D) were computed parcel-wise and averaged by hub. All paired $t$-tests are Bonferroni-corrected ($t(7)$, $p<0.001$).
The preceding analyses showed that the Early Visual, Ventral, and LOTC hubs form three distinct clusters embedded in two major processing streams. To uncover the content encoded in each hub, we decomposed the shared representational geometry using Kernel Multi-view Canonical Correlation Analysis (KMCCA)33. Standard RSA collapses two RDMs into a single Pearson’s $r$ that quantifies their alignment. By contrast, KMCCA treats each participant’s RDM as a separate “view” of the same stimulus-pair dissimilarities and identifies a shared subspace of orthogonal components whose projections are maximally correlated across all views. Our use of KMCCA parallels multi-view CCA techniques57–but instead of operating on raw voxel data, it works directly on the correlation kernels. These canonical axes therefore decompose the overall inter-subject alignment into ranked representational dimensions, isolating the latent components that drive shared geometry.
We focused our analysis on the first KMCCA component, which explains the largest fraction of inter-subject alignment. First, we identified the stimulus features associated with this component to interpret its semantic content (details in Supplementary Fig. A6). Second, we used partial RSA to quantify this component’s contribution to the overall alignment. In this analysis, a large reduction in RSA after controlling for the component would indicate a geometry dominated by that single axis, whereas a minimal change would imply a more complex, high-dimensional space with no single dominant factor.
The KMCCA projections revealed a distinct organization for each hub. In Early Visual Cortex (Figs. 5a and D8), stimuli formed a category-free cloud organized primarily by low-level visual features, consistent with its retinotopic properties rather than high-level semantics. In the Ventral hub (Figs. 5b and D9), the dominant axis traced a gradient from panoramic scenes to isolated objects, recapitulating the classic scene-to-object gradient in ventromedial cortex 27, 54. By contrast, the LOTC hub (Figs. 5c and D10) was organized along an animacy axis, with its first canonical component separating stimuli containing biological agents (people or animals) from inanimate scenes.
[…]
To determine whether this alignment drop in the LOTC hub was uniform across all images or driven specifically by biological content, we split the image set into stimuli containing animate agents (65.3%) and those without, then recomputed inter-subject alignment. The results confirmed the LOTC hub’s specialization: alignment vanished in the absence of biological elements in the scene, while it remained robust for scenes containing people or animals (Fig. 5e-f). An analysis of the connectivity network using this same data split showed that the entire lateral stream from the Early Visual Cortex to the LOTC hub was significantly reduced without biological content (Supplementary Fig. D7). Together, these findings provided support from a representational approach for a lateral “third visual pathway” specialized in social and animate information 30, 56.
Figure 6: Shared representational geometry is modulated by stimulus content and task. Alignment computed using the symmetric HCP atlas combining both hemispheres. (A–D) Replication in BOLD5000, using complex natural scenes. With a different task (valence rating) but with comparable complex scene stimuli that include social content, the main findings were replicated. Both the inter-subject alignment (A) and its correspondence with vision models (B) show a similar posterior-to-anterior progression. The layer-wise curves (C) reproduce the distinct hierarchical trend for each hub, and the connectivity matrix (D) reveals a similar network structure with a representationally distinct LOTC hub. (E–H) Replication in THINGS-fMRI, using object-centric images. With a stimulus set of isolated objects and a rapid oddball task, the extent of shared geometry is more restricted. Inter-subject alignment (E) and its correspondence with vision models (F) are largely confined to early visual cortex. While the layer-wise curves (G) retain their characteristic shapes, the connectivity matrix (H) is dominated by a single Early Visual hub. This suggests that rich scene and social content are necessary to engage the full processing hierarchy captured by shared representational geometry.
2.6 Generalization across tasks and stimulus sets
Our findings derived from NSD picture dataset showed that stimulus statistics can align brains with vision models during a long-format memory task, but representational geometry may depend on both image distribution and task context58, 59. To assess the degree to which our findings generalize to other encoding task environments, we replicated our analysis pipeline on two independent fMRI datasets: BOLD5000 (valence judgments on  5,000 natural scenes)24 and THINGS-fMRI (oddball detection on  9,000 object images)25. For each dataset, we derived single-trial voxel responses, computed inter-subject RSA maps, generated parcel-wise alignment scores for vision models, and computed a representational connectivity matrix using inter-subject, parcel-pair RSA (see Supplementary Section B for details).
To test the generalization of these findings, we first analyzed the BOLD5000 dataset 24. The stimulus set for BOLD5000 is highly comparable to NSD, containing complex natural scenes with multiple objects and social content, including a large subset of images drawn directly from the same database. Despite this stimulus similarity, participants performed a different task (valence judgment) with a shorter exposure time. Nevertheless, the results on images from BOLD500 dataset results largely replicated our main findings. Inter-subject and vision-model alignment spanned early visual cortex, ventral occipito-temporal parcels, and a LOTC hub (Fig. 6a,b), and the parcel-wise relationship between these two measures was well-described by the same power-law model ($R^{2}=0.70$). Furthermore, the layer-wise alignment curves recapitulated the same hierarchical trend observed in NSD: decreasing for the Early Visual Cortex, mixed for the Ventral Hub, and increasing for the LOTC Hub (Fig. 6c). The representational connectivity matrix also revealed a similar core network structure, with the LOTC hub forming a separate subsystem (Fig. 6d). These results indicate that the stimulus-driven brain-model convergence is robust to changes in task and scanning parameters.
Finally, we analyzed the THINGS-fMRI dataset 25, which presents a strong contrast in both stimulus content and task design. This dataset used briefly flashed (500ms), object-centric images that lack the complex scene and social content of NSD, and participants performed an orthogonal oddball task. We found that both inter-subject and vision-model alignment were largely confined to the early visual cortex (Fig. 6e,f), though the strong parcel-wise relationship between the two measures was preserved (power-law fit, $R^{2}=0.81$). Despite this restricted spatial extent, the underlying hierarchical profiles for each hub type were maintained: layer-wise curves still showed the characteristic decreasing, distributed, and increasing patterns (Fig. 6g). Consequently, the connectivity analysis revealed a single cluster in early visual cortex (Fig. 6h). This is consistent with a stimulus set lacking complex scene information and social interactions, and a passive task design that likely reduced the representational signal in higher-order cortical areas 55.
Across three distinct datasets, vision-model alignment was spatially co-located with inter-subject alignment, a finding consistent with models capturing the stimulus-driven geometry shared across brains. However, the extent of this processing cascade was context-dependent. Rich, complex scenes engaged the full posterior-to-LOTC hierarchy, whereas isolated objects under brief exposure primarily engaged only early visual cortex. These results therefore showed that shared representational geometry is a robust framework for revealing the specific cortical networks where stimulus information is processed across both individuals and models.

### v2

@@ 2.2 Model–brain alignment reveals modality-specific shared geometry
Figure 3: Hierarchical correspondence between model layers and cortical regions. (A–C) Layer-wise alignment (RSA) between vision models and representative parcels in the three hubs, averaged across all vision models. Panels show early visual cortex (A), the Ventral hub (B), and the LOTC hub (C). The $x$-axis expresses depth as a percentage of the total number of layers in each vision model. Curves show the mean alignment across participants ($N=8$); shaded bands indicate the standard error of the mean (SEM). Early visual parcels peak in shallow layers, Ventral parcels show distributed alignment across the hierarchy, and LOTC parcels peak in the deepest layers. (D) Cortical surface map of the vision-model layer that yields the highest alignment for each parcel (averaged across vision models). Colors encode normalized peak-layer depth (0% = shallowest layer, 100% = deepest layer), revealing a posterior–anterior gradient from shallow (green) to deep (purple) alignment. (E) Layer-wise alignment between language models and representative parcels from the three hubs (mean across language models). Alignment values are negligible in early visual and Ventral parcels and are restricted to LOTC, where curves rise quickly and then form a plateau rather than a smooth hierarchical progression. (F) Distribution across participants ($N=8$) of peak alignment depths for the 20 parcels with the highest overall vision–model alignment. For each subject and parcel, peak depth is defined as the mean depth (across vision models) of the layer at which RSA is maximal. Boxplots are ordered by median peak depth and colored by macro-anatomical cluster.
@@ 2.3 Hierarchical correspondence between models and cortex
To relate cortical parcels to the ordered representational stages learned by DNNs, we examined how model-brain alignment varied across model depth. Prior work has shown that, for vision models, shallow layers tend to align with early visual cortex and deeper layers tend to align with higher-level ventral and downstream regions 79, 51, 99, 17, 38. Here, we assessed this depth dependence during natural scene viewing in NSD by analyzing layer-wise RSA curves, focusing on the three hubs where inter-subject shared geometry was strongest (Fig. 3; full parcel-wise curves and model-family profiles in Supplementary Figs. 4-5).
The three hubs showed distinct depth profiles for vision models (Fig. 3a-c). Early visual parcels peaked in the shallowest layers and then decreased (Fig. 3a), consistent with correspondence to lower-level visual features. Ventral hub parcels showed broad alignment spanning earlier and later layers (Fig. 3b), indicating that their shared geometry relates to features distributed across multiple stages of the vision-model hierarchy rather than a single best-matching depth. LOTC parcels increased more steadily with depth and reached their maxima in deeper layers (Fig. 3c), indicating that the stimulus organization expressed in LOTC is best matched by later-stage vision-model features.
To summarize this organization across cortex, we identified for each parcel the vision-model layer with maximal alignment and projected this peak depth onto the cortical surface. The resulting map formed a posterior-to-anterior gradient in which early visual cortex was best explained by shallower layers, while more anterior occipitotemporal, STS-adjacent, and prefrontal regions were best explained by deeper layers (Fig. 3d). This gradient is consistent with hierarchical accounts of visual processing 28, 20, 54, 77 and with prior work showing that model depth tracks representational complexity across the human ventral pathway 38, 99, 98, 51, 17, 79. We use peak depth as a descriptive index that orders parcels along the model hierarchy (Fig. 3f).
Language models showed a different pattern. Alignment was concentrated in the LOTC hub, and layer-wise profiles did not show the early-to-late progression observed for vision models. Instead, alignment was present at the first layer, increased, and then approached a plateau. Across the language models we evaluated, first-layer alignment strongly predicted the plateau magnitude (Pearson’s $r=0.97$) and was related to the tokenizer representation (Pearson’s $r=0.78$; Supplementary Note 5). This pattern is consistent with evidence that, in visual contexts, brain-language correspondence can be driven by relatively coarse semantic content, closer to a bag-of-concepts organization than to progressively composed syntax 100. Interpreting this contrast depends on a linking assumption about inputs: vision models and cortex are compared while transforming pixel-level images, whereas language models operate on tokenized captions that already provide an abstracted description. Under this input mismatch, language-model alignment can emerge in regions that express higher-level scene organization, such as LOTC, without producing a depth-dependent gradient across the visual hierarchy.
@@
Figure 4: Inter-subject representational connectivity network. (A) Inter-subject connectivity matrix. Parcel-wise representational connectivity (RSA; Pearson’s $r$) for the 30 parcels with the highest mean IS-RSA ($N=8$). Each cell shows the mean RSA between the RDMs of two parcels, computed across all cross-subject pairs. The matrix shows three clusters corresponding to early visual cortex (V1–V4), the Ventral hub (VMV1–3, PHA1–3), and the LOTC hub (V4t, MT, MST, FST, TPOJ2–3). Warm colors indicate higher representational similarity; cool colors indicate dissimilarity. (B) Directed connectivity graph. The same parcels are shown as nodes. Node color encodes each parcel’s peak alignment depth with vision models (heuristic proxy for hierarchical position), node size indicates inter-subject alignment strength, and node border color indicates macro-anatomical group membership. Edges correspond to the union of three minimum spanning trees; edge color reflects the RSA value in panel A, and edge width indicates the spanning-tree iteration (first to third). Arrow direction is assigned from shallower to deeper peak-depth parcels when the depth difference is significant across subjects (paired $t$-test, $t(7)$, FDR-corrected $p<0.05$). The graph highlights two dominant routes from early visual cortex: a medial-ventral route toward the Ventral hub and a lateral route toward the LOTC hub.
2.4 A representational network linking ventromedial and lateral hubs
To characterize how regions with shared scene-evoked information relate to each other, we repurposed IS-RSA as a measure of representational connectivity. This defines a task- and stimulus-dependent network based on similarity in how stimuli are organized across regions, rather than on anatomical proximity or raw BOLD covariance. For each pair of parcels, we correlated their RDMs across different individuals, yielding a parcel-by-parcel connectivity matrix that quantifies how similarly distinct regions organize the stimulus set across observers (Fig. 4a). This approach emphasizes the informational content of the connectivity by grouping regions that share a common representational format, complementing univariate and conventional functional-connectivity analyses 69, 48.
The connectivity matrix showed three clusters aligned with the Early Visual, Ventral, and LOTC hubs (Fig. 4a). Early visual and ventral parcels formed a broad posterior cluster that also included the occipital and parietal cortex, indicating similar stimulus organization across these regions. LOTC parcels formed a second cluster with low or negative similarity to the posterior cluster. The strongest links between these systems passed through intermediate parcels in lateral occipital and ventral temporal cortex (LO1-3, FFC, PH). Because IS-RSA reflects both shared structure and measurement noise, local reliability differences can affect edge strength. However, the same topology was recovered with noise-ceiling normalization (Supplementary Fig. 6) and in a within-subject analysis (Supplementary Fig. 3).
To visualize the strongest routes, we extracted a sparse backbone graph by iteratively applying a minimum spanning tree algorithm to the parcel-wise connectivity matrix. We then oriented edges using each parcel’s peak alignment depth from the vision-model analysis (Fig. 3d,f) as a descriptive proxy for representational level, from shallower-layer peaks toward deeper-layer peaks. This ordering is not anatomical and does not imply causal direction. In the resulting graph, early visual cortex bifurcated into two routes (Fig. 4b): a medial-ventral path progressing through ventral occipitotemporal cortex into the Ventral hub, and a lateral path progressing through lateral occipital cortex into the LOTC hub. Extending the analysis to the whole brain (Supplementary Fig. 7) showed both routes continuing into association cortex, consistent with these representational formats interfacing with downstream systems implicated in memory, action, and control60, 75.
This two-route organization does not map cleanly onto the classical “ventral what / dorsal where” framing of visual processing 35, 28, 54, 90. The medial-ventral branch targeted parahippocampal and ventromedial visual parcels (including VMV and parahippocampal scene regions) that are consistently linked to scene layout and environmental context 27, 4, 26, and it matches proposals of a ventromedial, scene-based pathway that provides scene input to hippocampal memory systems 77, 78. The lateral route, by contrast, encompassed MT+ and LOTC and continued toward posterior STS and temporo-parietal cortex in the whole-brain graph, regions implicated in dynamic cues, biological motion, body perception, and social processing 6, 70, 78, 71, 15. In this representational network, regions typically associated with the classical dorsal stream in intraparietal and superior parietal cortex were less prominent in shared geometry, but they remained part of the broader posterior cluster. This pattern suggests that the LOTC-centered trajectory reflects a distinct lateral pathway whose shared stimulus organization is driven most strongly by biological and socially relevant content, rather than the spatial computations emphasized in traditional “where/how” accounts 70.
2.5 Representational dimensions driving shared geometry
Figure 5: KMCCA components and biological-content dependence of LOTC shared geometry. (A-C) Stimulus projections onto the first two KMCCA components for each hub, estimated from all eight NSD participants using the shared image set. Early Visual Cortex (A) shows no clear clustering by scene category. The Ventral hub (B) varies along a scene-object continuum. The LOTC hub (C) shows a categorical separation along an animacy-related axis (animate agents vs. non-animate scenes/objects). (D) Effect of removing the first KMCCA component (KMCCA1) using partial RSA. Orange boxplots show partial IS-RSA after regressing out KMCCA1 from the hub RDMs; blue shows the original IS-RSA. Removing this single dimension reduces inter-subject alignment in all hubs, with the largest reduction in LOTC. IS-RSA was computed per parcel and subject, averaged within hub, and then averaged across subjects; paired $t$-tests were Bonferroni-corrected ($t(7)$, $p<0.001$). Annotated $\Delta\rho$ reports the mean absolute change, and percentages denote the relative change normalized by baseline. (E-F) Inter-subject RSA maps recomputed after splitting the stimulus set into scenes with biological agents (E) versus without biological agents (F). LOTC alignment is strongly reduced when biological content is absent.
The two-route organization of the scene network raises a concrete question: which stimulus information most strongly organizes the shared representational geometry within each hub and along each route. To address this directly from the data, under the present task context, we decomposed shared across-subject structure in the Early Visual, Ventral, and LOTC hubs using Kernel Multi-view Canonical Correlation Analysis (KMCCA) 40. Unlike RSA, which summarizes similarity between two RDMs with a single correlation, KMCCA treats each participant’s hub RDM as a separate view of the same stimulus set and learns components that maximize correlation across views. We applied multi-view CCA ideas 86 to RDM-derived kernels, yielding a low-dimensional set of components that captures the geometry shared across participants within each hub.
To visualize this shared geometry, we projected stimuli onto the first two KMCCA components for each hub (Fig. 5a-c). In early visual cortex, the projection showed a diffuse distribution with no clear clustering by scene category, consistent with representations dominated by low-level image structure under this stimulus set and task 22 (Fig. 5a; Supplementary Fig. 9). In the Ventral hub, the leading axis formed a continuous gradient from panoramic scenes to more object-like images, consistent with ventromedial scene-object organization and scene layout/context representations 27, 77, 78, 4, 26 (Fig. 5b; Supplementary Fig. 10). In the LOTC hub, the first component separated images containing biological agents (people/animals) from a broader set of non-biological images, including objects and landscapes (Fig. 5c; Supplementary Fig. 11), consistent with LOTC sensitivity to biological and social content 70, 95, 6.
We then quantified how much this leading axis contributed to each hub’s shared geometry. We regressed out the first KMCCA component and recomputed inter-subject alignment using partial RSA (Fig.5d). Removing this single axis reduced IS-RSA in all three hubs (all $p<.001$, Bonferroni-corrected), but the reduction was substantially larger in LOTC ($30.4\%$) than in Early Visual Cortex ($8.0\%$; t(7)=10.78t(7)=-10.78, $p<.001$) or the Ventral hub ($3.7\%$; t(7)=12.21t(7)=-12.21, $p<.001$). This pattern indicates that an animacy-related dimension accounts for a large fraction of LOTC’s shared geometry, whereas Early Visual and Ventral hubs distribute shared variance across higher-dimensional structure (Supplementary Fig.20).
Given the dominant agent-related axis in LOTC, we tested whether lateral-route alignment depended on biological content by splitting the stimulus set into scenes with biological agents (people or animals; 65.3% of images) versus scenes without them, and recomputing IS-RSA within each subset. Alignment in LOTC was strongly reduced for non-biological scenes but remained robust when biological agents were present (Fig.5e–f). The same split also weakened the lateral branch of the representational network linking early visual cortex to LOTC, reducing the detectability of the two-route backbone when biological content was absent (Supplementary Fig. 8). This pattern indicates that biological-agent content is a major contributor to LOTC shared geometry in this stimulus set and task regime, consistent with proposals that lateral pathways are preferentially organized by socially and biologically relevant scene information.
Figure 6: Comparison across independent datasets. Analyses were performed on the BOLD5000 and THINGS-fMRI datasets using the same pipeline as NSD. (A–E) BOLD5000 (complex scenes, valence task; N=4). Despite lower signal reliability and differences in task, the core network topology persists in this dataset, which shares similar scene statistics with NSD. (A, B) Inter-subject alignment (A) and vision-model alignment (B) recover the characteristic occipitotemporal footprint, including Ventral and LOTC hubs, though with reduced magnitude and spatial continuity compared to NSD. (C) Layer-wise alignment profiles (mean $\pm$ SEM across subjects) preserve the hierarchical ordering (Early Visual << Ventral << LOTC) but exhibit flatter curves. (D, E) Representational connectivity matrix (D) and backbone graph (E) successfully recover the dual-stream topology (based on a $k=3$ minimum spanning tree), segregating the medial–ventral route from the lateral route. In the graph (E), node fill indicates peak model depth and edge color indicates representational connectivity strength. (F–J) THINGS-fMRI (object-centric images, oddball task; N=3). With minimal scene layout and social content, the network contracts. (F, G) Inter-subject (F) and vision-model (G) alignment are restricted primarily to early and ventral visual cortex, with no prominent LOTC peak. (H) Layer-wise profiles (mean $\pm$ SEM) are reduced in magnitude and show flatter profiles, especially in ventral and LOTC hubs. (I, J) The lateral stream does not emerge as a distinct subsystem in the connectivity matrix (I), and the backbone graph (J) lacks the clear two-route segregation evident in NSD and BOLD5000, consistent with the content-dependence of the lateral pathway.
2.6 Comparison across tasks and stimulus sets
Our analyses in NSD identified a distributed scene network by quantifying representational geometry shared across individuals and relating it to model feature spaces. Because this geometry is stimulus- and task-dependent, its spatial expression can vary with the regularities present in the stimulus set 19, 41, task demands that reweight the stimulus features that are emphasized 8, 45, and measurement reliability, which bounds how well shared structure can be estimated from trial-level data 63, 21, 92. To assess how the inferred organization generalizes across experimental regimes, we applied the same pipeline to two independent datasets: BOLD5000 11 and THINGS-fMRI 46 (Fig.6).
BOLD5000 11 uses natural images broadly similar to NSD but differs in task (valence judgment vs. memory), stimulus duration (1 s vs. 3 s), and trial-level reliability under comparable estimation procedures 2. In this regime, inter-subject and vision-model alignment remained concentrated in occipital and occipitotemporal cortex, with clear involvement of the ventral and LOTC hubs identified in NSD (Fig. 6a-b), consistent with both datasets sampling rich scene layout, context, and social content. However, the overall spatial profile differed from NSD in two main respects. Early visual cortex was not detected as a dominant peak, and the LOTC pattern was less spatially contiguous. These differences are likely attributable to reduced signal reliability resulting from the lower field strength (3T) and shorter presentation times. Despite these differences, vision-model alignment preserved a monotonic parcel-wise relationship with IS-RSA, as in NSD (power-law scaling $R^{2}=0.71$). Layer-wise profiles preserved the hub-specific ordering but were flatter than in NSD (Fig. 6c).
At the network level, representational connectivity in BOLD5000 still separated a broad posterior cluster spanning early visual, ventral, and adjacent cortex from a lateral occipitotemporal cluster that included LOTC together with nearby lateral occipital and ventral temporal parcels (Fig. 6d). LOTC was less segregated from neighboring lateral parcels than in NSD, but the backbone visualization still recovered a two-route organization from early visual cortex toward ventral versus lateral targets (Fig. 6e), consistent with a ventromedial route associated with scene context and a lateral route associated with biological and social features.
THINGS-fMRI differs from NSD in both stimulus content and task. It uses briefly presented (500 ms), object-centric images with minimal scene layout and reduced social content, and participants performed an orthogonal oddball task 46. In this regime, inter-subject alignment and vision-model alignment again remained tightly coupled (power-law scaling $R^{2}=0.80$; Fig. 6f-g), but the spatial pattern was concentrated in occipital and ventral visual cortex. The NSD ventromedial/parahippocampal and LOTC hubs were not detected as prominent hubs, and LOTC resembled the NSD pattern obtained when restricting analyses to images without biological agents (Fig. 5f). Layer-wise profiles retained the same ordering but were reduced in magnitude and less distinctive for the ventral and LOTC hubs (Fig. 6h). Consistent with this, representational connectivity did not separate an LOTC-centered subsystem (Fig. 6i), and the backbone graph did not recover the two-route topology observed for NSD and BOLD5000 (Fig. 6j).
Across datasets, vision-model alignment continued to track inter-subject alignment, indicating that the same stimulus-locked geometry that is detectable across observers is also present in image-trained feature spaces. However, which hubs and routes emerged as prominent depended on the structure available in the stimulus set and on task/reliability constraints: scene-rich datasets expressed ventromedial/parahippocampal and LOTC involvement more clearly, whereas the object-centric THINGS regime emphasized early and ventral visual cortex and did not express a separable LOTC-centered route.